How much potential energy does a car gain if a crane lifts the car with a mass of 1,500 kg and 20 m straight up


1
Expert's answer
2021-12-02T14:07:19-0500

"PE=m\times g \times h"

where "PE=" Potential energy

"m=" Mass of the car

"g=" Gravity

"h=" Height

Therefore:

"PE= 1500kg\u00d79.81M\/s^2\u00d720M=294.3kJ"